OpenStudy (fouzberzerk):

Will fan and medal. -5/4n = 40 Find N's value

rvc (rvc):

cross multiply

OpenStudy (fouzberzerk):

?

rvc (rvc):

-5=40 x 4n

OpenStudy (fouzberzerk):

idk.

OpenStudy (fouzberzerk):

I'm just being honest with you here, I don't know.

rvc (rvc):

-5/160=n

OpenStudy (fouzberzerk):

-0.03125

OpenStudy (fouzberzerk):

Can I just get an answer here? I've been spending my entire night on one problem.

rvc (rvc):

thats the answer

OpenStudy (fouzberzerk):

No

OpenStudy (fouzberzerk):

Finally, someone new.

OpenStudy (fouzberzerk):

Help?

OpenStudy (er.mohd.amir):

-5=4n*40 160n=-5 n=-5/160 n=-0.03125

OpenStudy (fouzberzerk):

That's not going to work.

OpenStudy (fouzberzerk):

It only lets me put a few numbers in.

OpenStudy (fouzberzerk):

4 characters to be exact.

OpenStudy (er.mohd.amir):

what the problem

OpenStudy (fouzberzerk):

Maybe I'm typing the problem wrong. Here's the actual problem.

OpenStudy (fouzberzerk):

http://imgur.com/YNi9D4z

OpenStudy (er.mohd.amir):

that is -5n/4=40 -5n=40*4 -n=40*4/5 n=-32

OpenStudy (fouzberzerk):

Great, I need help with another question now.

OpenStudy (fouzberzerk):

http://imgur.com/FM6diWe

OpenStudy (er.mohd.amir):

-17/7+b=43/7 b=43/7+17/7 b=60/7 b=8 4/7

OpenStudy (fouzberzerk):

Alright, one or two more.

OpenStudy (fouzberzerk):

http://i.imgur.com/qfrPcSu.png?1

OpenStudy (er.mohd.amir):

1/4+w=19/3 w=19/3-1/4 w=(19*4-1*3)/12 w=73/12=6 1/12

OpenStudy (fouzberzerk):

It has to equal 6 1/3

OpenStudy (er.mohd.amir):

no 6 and 1/12

OpenStudy (fouzberzerk):

6 and 1/3 look at the problem.

OpenStudy (fouzberzerk):

1/4 + w = 6 1/3

OpenStudy (er.mohd.amir):

6 1/4=19/3

OpenStudy (fouzberzerk):

1/3

OpenStudy (er.mohd.amir):

w=19/3-1/4 tranfer the 1/4

OpenStudy (fouzberzerk):

That comes out to 6 7/12.

OpenStudy (er.mohd.amir):

take LCM and solve it w=19/3-1/4

OpenStudy (fouzberzerk):

?

OpenStudy (er.mohd.amir):

(19*4-1*3)/12

OpenStudy (fouzberzerk):

73/12?

OpenStudy (er.mohd.amir):

correct

OpenStudy (fouzberzerk):

6 1/2 simplified?

OpenStudy (er.mohd.amir):

73 divide by 12 get 6 and remainder is 1 so 73/12=6 1/12

OpenStudy (fouzberzerk):

That still doesn't solve my question. I need something that equals 6 1/3

OpenStudy (fouzberzerk):

Whatever, forget about that question.

OpenStudy (fouzberzerk):

Help me with this one: http://imgur.com/RIxKSmj

OpenStudy (er.mohd.amir):

subtract tallen and smaller

OpenStudy (fouzberzerk):

1 5/6?

OpenStudy (er.mohd.amir):

19/3-9/2

OpenStudy (fouzberzerk):

What?

OpenStudy (er.mohd.amir):

6 1/3=(6*3+1)/3=19/3 and 4 1/2=(4*2+1)/2=9/2

OpenStudy (fouzberzerk):

What's the final answer dude. I'm extremely burnt out and I just want some sleep.

OpenStudy (er.mohd.amir):

19/3-9/2=(19*2-9*3)/6 (38-27)/6=11/6 1 5/6

OpenStudy (fouzberzerk):

1 more.

OpenStudy (er.mohd.amir):

B

OpenStudy (fouzberzerk):

http://i.imgur.com/DMR7SHD.png?1

OpenStudy (fouzberzerk):

Just a direct answer again, I need sleep.

OpenStudy (er.mohd.amir):

x=6-1/6=(36-1)/*6=35/6 =5 5/6 that is C

OpenStudy (er.mohd.amir):

here at 12:30 PM that is noon

OpenStudy (er.mohd.amir):

and U going to sleep

OpenStudy (fouzberzerk):

2:45 AM here, I'm preparing for a big game release on Monday. Getting all my missing work done.

OpenStudy (er.mohd.amir):

this is because of internet

OpenStudy (fouzberzerk):

Let's see what my final precentage is.

OpenStudy (fouzberzerk):

I got an 80% which is a B, thanks alot. And yeah.